Multiply.
(s+(1)/(4))(s+(1)/(5))
Is this the correct answer?: s^2+9/20s+1
Anyone please help me. Thank you.

Answer by jorel1380(3719)

(s+1/4)(s+1/5)=s2+s/4+s/5+1/20=s2+9/20+1/20..